OpenStudy (loser66):

If A is linear transformation of rank one, then there exists a unique scalar \(\alpha\) such that \(A^2= \alpha A\) b) If \(\alpha\neq1\), then 1- A is invertible Please, help

OpenStudy (loser66):

Please, I do not know how to start. I have a formula: Ax = [x,y0]x0 but I don't know how to use it. hihihi... stupid me!!

OpenStudy (kinggeorge):

Well, if it has rank 1, then all the rows are scalar multiples of the first row. So if we apply the linear transformation to a vector \((x_1,...,x_n)^T\), then we get\[A(x_1,...,x_n)^T=(a,a_2\cdot a,...,a_n\cdot a)^T\]for some constants \(a_2,...,a_n\). Applying the transformation again, we get\[A(a,a_2\cdot a,...,a_n\cdot a)^T=(b,a_2\cdot b,...,a_n\cdot b)^T.\]Let \(b=\alpha a\). Then we have the relations\[A^2(x_1,...,x_n)^T=(\alpha a,a_2\cdot \alpha a,...,a_n\cdot \alpha a)^T=\alpha(a,a_2\cdot a,...,a_n\cdot a)^T.\]This last bit is exactly \(A(x_1,...,x_n)^T\), so \(A^2=\alpha A\) for some constant \(\alpha\).

OpenStudy (kinggeorge):

Typo. The last line should read: This last bit is exactly \(\alpha A(x_1,...,x_n)^T\), so \(A^2=\alpha A\) for some constant \(\alpha\).

OpenStudy (loser66):

I am sorry, I don't get the first equation A is linear transformation, let say defined by [A] = (a_ij) how you get the right hand side?

OpenStudy (kinggeorge):

Let \(a_{11},...,a_{1n}\) be the first row of the matrix. Then we can say that \[a_{11}x_1+...+a_{1n}x_n=a.\]Since the rows are all multiples of the first row, let the \(i\)-th row be \(a_i\) times the first row. So the \(i\)-th row is \(a_ia_{11},...,a_ia_{1n}\). Thus, \[a_ia_{11}x_1+...+a_ia_{1n}x_n=a_ia\]

OpenStudy (loser66):

Thanks so much, I got it now. Please, give me a hint for b)

OpenStudy (kinggeorge):

I'm not actually sure about part b myself right now.

OpenStudy (loser66):

No worry, I still have time for this problem. :)

OpenStudy (kinggeorge):

Have you done anything with eigenvalues/eigenvectors?

OpenStudy (kinggeorge):

In any case, since \(A^2=\alpha A\), it must be that \(A\vec{x}=\alpha\vec{x}\) for any vector \(\vec{x}\). Also, we know that \(I-A\) is NOT invertible if and only if \(\det(I-A)=0\) if and only if there is some non-zero vector \(\vec{x}\) such that \[\vec{0}=(I-A)\vec{x}=I\vec{x}-A\vec{x}=\vec{x}-\alpha\vec{x}=(1-\alpha)\vec{x}.\]Since \(\vec{x}\) is non-zero, the rightmost term is zero if and only if \(\alpha=1\). So it follows that \(I-A\) is not invertible if and only if \(\alpha=1\), and therefore if \(\alpha\neq1\), then \(I-A\) must be invertible.